Analysis and Interpretation of Profitability

Balance sheets and income statements for 3M Company follow.

Consolidated Statements of Income
Years ended December 31 ($ millions) 2009 2008 2007
Net sales $23,123 $25,269 $24,462
Operating expenses
Cost of sales 12,109 13,379 12,735
Selling, general and administrative expenses 4,907 5,245 5,015
Research, development and related expenses 1,293 1,404 1,368
Loss/(gain) from sale of business -- 23 (849)
Total operating expenses 18,309 20,051 18,269
Operating income 4,814 5,218 6,193
Interest expenses and income
Interest expense 219 215 210
Interest income (37) (105) (132)
Total interest expense 182 110 78
Income before income taxes 4,632 5,108 6,115
Provision for income taxes 1,388 1,588 1,964
Net income including noncontrolling interest 3,244 3,520 4,151
Less: Net income attributable to noncontrolling interest 51 60 55
Net income $ 3,193 $ 3,460 $ 4,096
Consolidated Balance Sheets
($ millions) 2009 2008
Assets
Current Assets
Cash and cash equivalents $ 3,040 $ 1,849
Marketable securities-current 744 373
Accounts receivable-net 3,250 3,195
Inventories
Finished goods 1,255 1,505
Work in process 815 851
Raw materials and supplies 569 657
Total inventories 2,639 3,013
Other current assets 1,122 1,168
Total current assets 10,795 9,598
Marketable securities-noncurrent 825 352
Investments 103 111
Property, plant and equipment 19,440 18,812
Less: Accumulated depreciation (12,440) (11,926)
Property, plant and equipment-net 7,000 6,886
Goodwill 5,832 5,753
Intangible assets-net 1,342 1,398
Prepaid pension benefits 78 36
Other assets 1,275 1,659
Total assets $ 27,250 $ 25,793
Liabilities
Current liabilities
Short-term borrowings and current portion of long-term debt $ 613 $ 1,552
Accounts payable 1,453 1,301
Accrued payroll 680 644
Accrued income taxes 252 350
Other current liabilities 1,899 1,992
Total current liabilities 4,897 5,839
Long-term debt 5,097 5,166
Pension and postretirement benefits 2,227 2,847
Other liabilities 1,727 1,637
Total liabilities 13,948 15,489
Equity
3M Company shareholders' equity: Common stock, par value $.01 per share; 9 9
Additional paid-in capital 3,153 3,006
Retained earnings 23,753 22,227
Treasury stock (10,397) (11,676)
Accumulated other comprehensive income (loss) (3,754) (3,686)
Total 3M Company shareholders' equity 12,764 9,880
Noncontrolling interest 538 424
Total equity 13,302 10,304
Total liabilities and equity $ 27,250 $ 25,793

(a) Compute net operating profit after tax (NOPAT) for 2009. Assume that the combined federal and statutory rate is: 36.1% (Round your answer to the nearest whole number.)
2009 NOPAT = ? ($ millions)

(b) Compute net operating assets (NOA) for 2009 and 2008. Treat noncurrent Investments as a nonoperating item.
2009 NOA = ? ($ millions)
2008 NOA = ? ($ millions)

(c) Compute 3M's RNOA, net operating profit margin (NOPM) and net operating asset turnover (NOAT) for 2009. (Round your answers to two decimal places. Do not use NOPM x NOAT to calculate RNOA.)
2009 RNOA = ? %
2009 NOPM = ? %
2009 NOAT = ?

(d) Compute net nonoperating obligations (NNO) for 2009 and 2008.
2009 NNO = ? ($ millions)
2008 NNO = ? ($ millions)

(e) Compute return on equity (ROE) for 2009. (Round your answers to two decimal places. Do not round until your final answer.)
2009 ROE = ? %

(f) What is the nonoperating return component of ROE for 2009? (Round your answers to two decimal places.)
Hint: Use your prior rounded answers to compute this answer.
2009 nonoperating return = ? %
Answer #1

Part A

2009 NOPAT = Operating income -Tax on operating profit

Tax on operating profit = Tax Expense + Tax Shield= Tax Expense + (Pretax net nonoperating expense × Statutory tax rate)

=4814-(1388+(182*36.1%)) = $3360

Part B

NOA = Operating Assets – Operating Liabilities

NOA = (total assets – cash and equivalents – Marketable securities-current Marketable securities-noncurrent – investments) – (Total liabilities - Short-term borrowings and current portion of long-term debt - Long-term debt)

2009 NOA = (27250-3040-744-825-103)-(13948-163-5097) = $13850

2008 NOA = (25793-1849-373-352-111)-(15489-1552-5166) = $14337

Part C

2009 RNOA = NOPAT/Average NOA = 3360/((13850+14337)/2) = 23.84%

2009 NOPM = NOPAT/Sales = 3360/23123 = 14.53%

2009 NOAT = Sales/ Average NOA = 23123/((13850+14337)/2) = 1.64

Part D

NNO = Nonoperating Liabilities + Noncontrolling Interest – Nonoperating Assets

2009 NNO = (163+5097) + 0 – (3040+744+825+103) = $548

2008 NNO = (1552+5166) +0 – (1849+373+352+111) = 4033

Part E

ROE = NI/Average SE = 3193/((12764+9880)/2)= 28.20%

Part F

Nonoperating Return = ROE – RNOA = 28.20%-23.84% = 4.36%

Part G

Option B

ROE > RNOA implies that 3M is able to borrow money to fund operating assets that yield a return greater than its cost of debt.